A chemist needs 150 milliliters of a 44% solution but has only 24% and 99% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
Answer by ankor@dixie-net.com(22740)

A chemist needs 150 milliliters of a 44% solution but has only 24% and 99% solutions available.
Find how many milliliters of each that should be mixed to get the desired solution.
:
Let x = amt of 99% solution required
the resulting amt is to be 150 ml, therefore:
(150-x) = amt of 24% solution
:
A typical mixture equation
:
.99x + .24(150-x) = .44(150)
.99x + 36 - .24x = 66
.99x - .24x = 66 - 36
.75x = 20
x = 30/.75
x = 40 ml of the 99% solution
then
150 - 40 = 110 ml of the 24% solution
:
:
You should check this in the original mixture equation
.99(40) + .24(110) = .44(150)